# Apparent Phantom Crossing in Gauss-Bonnet Gravity

### v1

@@ I Introduction
The accelerated expansion of the present Universe could be generated by an unknown fluid filling the Universe, which is dubbed dark energy. If the dark energy is a perfect fluid, it is characterized by the ratio $w$ of the pressure $p$ and the energy density $\rho$ of the dark energy fluid, $w=\frac{p}{\rho}$, which is called the equation of state (EoS) parameter. The cosmological constant corresponds to w=1w=-1. The dark energy with $\frac{1}{3}>w>-1$ is called quintessence and that with w<1w<-1 is phantom. The quintessence dark energy can be realized by a canonical scalar field with a potential, which could be an analogy with the inflaton, which generates the inflation in the early Universe. Recently, the observations of the Dark Energy Spectroscopic Instrument (DESI) [1, 2] seem to indicate that there could have been a transition from the phantom Universe to the non-phantom Universe. The transition is called the phantom crossing [5, 6, 7, 8]. For a mainstream of articles related to the DESI data, see Refs. [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. Formally, the phantom crossing means the transition from the non-phantom Universe to the phantom Universe. Therefore, we may call this transition inverse phantom crossing. The observation of DESI seems to indicate that the EoS parameter $w$ could change depending on time. Furthermore, the phantom dark energy cannot be realized by the canonical scalar field without resorting to tachyon fields. Therefore, we need a model of dark energy distinct from the canonical scalar field framework.
@@
which gives the time-dependence of the coefficient function $\xi(\phi)$. We should note, however, that the condition (5) cannot be satisfied in a general spacetime, like a black hole spacetime [48], except for a trivial case that $\xi(\phi)$ is a constant. We should also note that the observed oldest event of the merger of a black hole and a neutron star occurred when the redshift $z\sim 0.5$, which could be 5-8 billion years ago depending on models [4]. Therefore, there is no constraint like (4) of the time-dependence of $\xi(\phi)$ as in (5) before one billion years ago. For example, the (inverse) phantom crossing observed in the DESI observation could have occurred about 5-8 billion years ago. It is possible that the coefficient function $\xi(\phi)$ may have drastically evolved and played a dominant role in the early Universe.
@@
Although the phenomenon of the phantom crossing is still not established, and it might be an artifact of the choice in parametrization, in this paper, we construct realistic models showing the inverse phantom crossing in the framework of the scalar–Einstein–Gauss-Bonnet gravity and the ghost-free $f(G)$ gravity. We also construct models where the standard phantom crossing and the inverse phantom crossing iterate alternately. Recently, in [59], a new scenario, which might solve the problem in the DESI observations, was proposed. In this scenario, if the energy density of the dark matter decreases more slowly than $a^{-3}$, the DESI observations might be explained without the real transition from w<1w<-1 to w>1w>-1 of the EoS parameter of dark energy. We consider this scenario by using the scalar–Einstein–Gauss-Bonnet gravity and the ghost-free $f(G)$ gravity. In this scenario, the total energy density and the total pressure do not violate the energy conditions; the components included in the total energy and pressure can satisfy the energy conditions. Furthermore, we consider the scenario by assuming that the particles corresponding to the scalar field $\phi$ in the action (1) become dark matter. Due to the coupling $\xi(\phi)$ with the Gauss-Bonnet invariant $\mathcal{G}$ in (1), the mass depends on the curvature and by properly choosing the potential $V(\phi)$ and the coupling $\xi(\phi)$, the energy density can decrease slower than $a^{-3}$. This last scenario also suggests that the inverse phantom crossing might occur due to the transition from the decelerating expansion of the Universe to the accelerating expansion.
II Scalar–Einsten–Gauss-Bonnet gravity
@@
$0=$ $\,-\ddot{\phi}-3H\dot{\phi}-V^{\prime}(\phi)-24H^{2}\left(H^{2}+\dot{H}\right)\,$, (12)
@@ IV Phantom Crossing in Ghost-free $f(\mathcal{G})$ Gravity
In the scalar–Einstein–Gauss-Bonnet gravity (1), the propagation of the scalar field $\phi$ might generate the fifth force, and the fluctuation may affect the structure formation in the Universe. In order to avoid such problems, an easy way could be to eliminate the scalar field. The $f(\mathcal{G})$ gravity modell [50, 51, 52, 53, 54, 55] was one of such candidates.
@@ V.3 Apparent phantom crossing and energy conditions
which indicates that $w_{\mathcal{G}}<-1$, when $N>N_{0}$ but $w_{\mathcal{G}}<-1$, when $N<N_{0}$. Therefore, all the energy conditions are violated when $N<N_{0}$.
@@ V.4 Scalar particle scenario
$0=$ $\,-\frac{3}{\kappa^{2}}H^{2}+\frac{1}{2}H^{2}{\phi^{\prime}}^{2}+V_{0}+\frac{1}{2}{m_{0}}^{2}\phi^{2}+24H^{4}{m_{1}}^{2}\phi\phi^{\prime}+\rho_{\phi\mathrm{DM}\,0}\sqrt{{m_{0}}^{2}+24{m_{1}}^{2}H^{2}\left(H^{2}+\dot{H}\right)}\mathrm{e}^{-3N}+\sum_{m}\rho_{m0}\mathrm{e}^{-3\left(1+w_{m}\right)N}\,.$ (73)
@@ VI Summary and Discussions
Acknowledgements
This work was partially supported by the program Unidad de Excelencia Maria de Maeztu CEX2020-001058-M, Spain (S.D. Odintsov). This research has been funded by the Committee of Science of the Ministry of Education and Science of the Republic of Kazakhstan (Grant No. AP26194585) (S.D. Odintsov and V.K. Oikonomou).

### v2

@@ I Introduction
The accelerated expansion of the present Universe could be generated by an unknown fluid filling the Universe, which is dubbed dark energy. If the dark energy is a perfect fluid, it is characterized by the ratio $w$ of the pressure $p$ and the energy density $\rho$ of the dark energy fluid, $w=\frac{p}{\rho}$, which is called the equation of state (EoS) parameter. The cosmological constant corresponds to w=1w=-1. The dark energy with $\frac{1}{3}>w>-1$ is called quintessence and that with w<1w<-1 is phantom. The quintessence dark energy can be realized by a canonical scalar field with a potential, which could be an analogy with the inflaton, which generates the inflation in the early Universe. Recently, the observations of the Dark Energy Spectroscopic Instrument (DESI) [1, 2] seem to indicate that there could have been a transition from the phantom Universe to the non-phantom Universe. The transition is called the phantom crossing [5, 6, 7, 8]. For a mainstream of articles related to the DESI data, see Refs. [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. Formally, the phantom crossing means the transition from the non-phantom Universe to the phantom Universe. Therefore, we may call this transition inverse phantom crossing. The observation of DESI seems to indicate that the EoS parameter $w$ could change depending on time. Furthermore, the phantom dark energy cannot be realized by the canonical scalar field without resorting to tachyon fields. Therefore, we need a model of dark energy distinct from the canonical scalar field framework.
@@
which gives the time-dependence of the coefficient function $\xi(\phi)$. We should note, however, that the condition (5) cannot be satisfied in a general spacetime, like a black hole spacetime [50], except for a trivial case that $\xi(\phi)$ is a constant. We should also note that the observed oldest event of the merger of a black hole and a neutron star occurred when the redshift $z\sim 0.04$, which could be about one billion years ago depending on models [4]. Therefore, there is no constraint like (4) of the time-dependence of $\xi(\phi)$ as in (5) before one billion years ago. For example, the (inverse) phantom crossing observed in the DESI observation could have occurred about 5-8 billion years ago corresponding to $z\sim 0.5$. It is possible that the coefficient function $\xi(\phi)$ may have drastically evolved and played a dominant role in the early Universe.
@@
Although the phenomenon of the phantom crossing is still not established, and it might be an artifact of the choice in parametrization, in this paper, we construct realistic models showing the inverse phantom crossing in the framework of the scalar–Einstein–Gauss-Bonnet gravity and the ghost-free $f(G)$ gravity [^1] For the model of the dark energy in the framework of the scalar–Einstein–Gauss-Bonnet gravity from the point view of particle physics, see [61, 62] although the phantom crossing was not found there. . We also construct models where the standard phantom crossing and the inverse phantom crossing iterate alternately. Recently, in [63], a new scenario, which might solve the problem in the DESI observations, was proposed. In this scenario, if the energy density of the dark matter decreases more slowly than $a^{-3}$, the DESI observations might be explained without the real transition from w<1w<-1 to w>1w>-1 of the EoS parameter of dark energy. We consider this scenario by using the scalar–Einstein–Gauss-Bonnet gravity and the ghost-free $f(G)$ gravity. In this scenario, the total energy density and the total pressure do not violate the energy conditions; the components included in the total energy and pressure can satisfy the energy conditions. Furthermore, we consider the scenario by assuming that the particles corresponding to the scalar field $\phi$ in the action (1) become dark matter. Due to the coupling $\xi(\phi)$ with the Gauss-Bonnet invariant $\mathcal{G}$ in (1), the mass depends on the curvature and by properly choosing the potential $V(\phi)$ and the coupling $\xi(\phi)$, the energy density can decrease slower than $a^{-3}$. This last scenario also suggests that the inverse phantom crossing might occur due to the transition from the decelerating expansion of the Universe to the accelerating expansion.
II Scalar–Einstein–Gauss-Bonnet gravity
@@
$0=$ $\,-\ddot{\phi}-3H\dot{\phi}-V^{\prime}(\phi)-24H^{2}\left(H^{2}+\dot{H}\right){\xi^{\prime}(\phi)}\,$, (12)
@@ III.1 A Realistic Model
Eq. (25) tells that when $\alpha\left(N-N_{0}\right)\to\pm\infty$, $\rho_{\mathcal{G}}\to\Lambda_{0}$. THis tells $V(\phi)\to\Lambda_{0}$ and $\xi(\phi)\to\mathrm{constant}$ when $\alpha\left(N-N_{0}\right)\to\pm\infty$ in the action (1) and therefore the constraint (4) coming from the gravitational wave could be satisfied in the late universe. As we mentioned, the observed oldest event of the merger of a black hole and a neutron star corresponds to the redshift $z\sim 0.04$. On the other hand, the phantom crossing observed DESI corresponds to $z\sim 0.5$ and therefore $\mathrm{e}^{N_{0}}=\frac{1}{1+0.5}\sim 0.67$. Let the number of the $e$-foldings corresponding to the oldest event of the merger of a black hole and a neutron star be $N_{\mathrm{mrgr}}$. Then we find $\mathrm{e}^{N_{\mathrm{mrgr}}}=\frac{1}{1+0.04}\sim 0.96$ and we find $\mathrm{e}^{\alpha\left(N_{\mathrm{mrgr}}-N_{0}\right)}\sim\left(1.44\right)^{\alpha}$. Therefore, if we choose $\alpha$ large enough, the constraint (4) could be satisfied
[…]
@@ IV Phantom Crossing in Ghost-free $f(\mathcal{G})$ Gravity
In the scalar–Einstein–Gauss-Bonnet gravity (1), the propagation of the scalar field $\phi$ might generate the fifth force, and the fluctuation may affect the structure formation in the Universe. In order to avoid such problems, an easy way could be to eliminate the scalar field. The $f(\mathcal{G})$ gravity model [52, 53, 54, 55, 56, 57] was one of such candidates.
@@ V.1 Slowly decreasing energy density of dark matter
You may suspect that the conservation law for the scalar particle might not be violated. This is because, due to the Gauss-Bonnet coupling, there is an exchange of energy with the geometry.
@@ V.3 Apparent phantom crossing and energy conditions
which indicates that $w_{\mathcal{G}}>-1$, when $N>N_{0}$ but $w_{\mathcal{G}}<-1$, when $N<N_{0}$. Therefore, all the energy conditions are violated when $N<N_{0}$.
@@ V.4 Scalar particle scenario
$0=$ $\,-\frac{3}{\kappa^{2}}H^{2}+\frac{1}{2}H^{2}{\phi^{\prime}}^{2}+V_{0}+\frac{1}{2}{m_{0}}^{2}\phi^{2}+24H^{4}{m_{1}}^{2}\phi\phi^{\prime}+\rho_{\phi\mathrm{DM}\,0}\sqrt{{m_{0}}^{2}+24{m_{1}}^{2}H^{2}\left(H^{2}+\dot{H}\right)}\mathrm{e}^{-3N}$
$\,+\sum_{m}\rho_{m0}\mathrm{e}^{-3\left(1+w_{m}\right)N}\,.$ (73)